You are hoping to buy a house in the future and recently received an inheritance of

​$24,000.

You intend to use your inheritance as a down payment on your house.

a.  If you put your inheritance in an account that earns 9 percent interest compounded​ annually, how many years will it be before your inheritance grows to $31,000​?

b.  If you let your money grow for 10.5 years at 9 percent​, how much will you​ have?

c.  How long will it take your money to grow to ​$31,000 if you move it into an account that pays 5 percent compounded​ annually? How long will it take your money to grow to  $31,000 if you move it into an account that pays 11 percent​?

d.  What does all this tell you about the relationship among interest​ rates, time, and future​ sums?
Answer #1

PV =24000
a. Rate =9%
FV =31000
FV =PV*(1+r)^n
31000=24000*(1+9%)^n
Applying log on both sides we get
N =log(31000/24000)/log(1.09) =2.97 years

b. Rate =9%
Number of Years =10.5
FV =PV*(1+r)^n =24000*(1+9%)^10.5 =59318.41

c. Rate =5%
PV =24000
FV =31000
FV =PV*(1+r)^n
31000=24000*(1+5%)^n
Applying log on both sides we get
N =log(31000/24000)/log(1.05) =5.25 years

If rate =11%
Applying log on both sides we get
N =log(31000/24000)/log(1.11) =2.45 years

d. At higher interest rate time required will be less to achieve same FV.
If interest rate and time is high then FV achieved for same amount is high.